The velocity v of a wave in a stretched string depends on tension T, in the string and the mass per unit length U of the string.Obtain an expression for V in terms of T and U, using the method of dimensions


1
Expert's answer
2021-09-15T11:38:17-0400

Explanations & Calculations


  • We can write the relationship between the quantities as follows

"\qquad\qquad
\begin{aligned}
\small v&\propto \small T^m.U^n
\end{aligned}"

  • Then comparing both sides for their dimensions, m and n can be found which helps writing a tentative relationship.

"\qquad\qquad
\begin{aligned}
\small [v]&=\small LT^{-1}\\
\small [T]&=\small MLT^{-2}\\
\small [U]&=\small \frac{[mass]}{[length]}=\frac{M}{L}=ML^{-1}\\
\\
\small L.H.S &: \small LT^{-1}\\
\small R.H.S &:\small (MLT^{-2})^m.(ML^{-1})^n\\
&:\small M^{(m+n)}.L^{(m-n)}.T^{(-2m)}\\
\\
\small index(L.H.S) &: \small index(R.H.S)\\
\small M \to 0&=\small m+n \implies m=-n\\
\small L \to 1&=\small m-n \\
&\implies \small m=\frac{1}{2},n=\frac{-1}{2}\\
\small T \to -2&=\small -2m\,(proves\,the\,results)\\
\\
\therefore\, v&\propto \small T^{\frac{1}{2}}.U^{\frac{-1}{2}}\\
&\propto\small \sqrt{\frac{T}{U}}
\end{aligned}"

  • Then a acceptable equation would be

"\qquad\qquad
\begin{aligned}
\small v&=\small k.\sqrt{\frac{T}{U}}
\end{aligned}"

  • Value of k to be found experimentally.